1. What is the molarity of a solution which contains 5.85 g of NaCl in 100 L of water?
2. Which of the following processes increases the rate at which most ionic solids dissolve in water?
Agitation
Large surface area
decreasing the temperature
Agitation and larger surface area
1
Expert's answer
2020-05-04T01:52:28-0400

Solution.

1.

"m(NaCl)=5.85g;"

"V(H_2O)=100L;"

"M(NaCl)=58g\/mol;"

"Molarity=\dfrac{mol solute}{L solution};"

"\nu(NaCl)=\dfrac{m(NaCl)}{M(NaCl)};"

"\nu(NaCl)=\dfrac{5.85g}{58g/mol}=0.1mol;"

"Molarity=\dfrac{0.1}{100}=0.0001mol/L;"

2.Agitation and larger surface area;

Answer: 1. "Molarity=0.0001mol\/L;"

2.Agitation and larger surface area.